Guile-GNOME: GNOME-VFS

version 2.16.2, updated 9 December 2011

The Free Software Foundation Christophe Fergeau Christian Neumair

This manual is for (gnome vfs) (version 2.16.2, updated 9 December 2011) Copyright 2001-2007 The Free Software Foundation, Christophe Fergeau, Christian Neumair

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.1 or any later version published by the Free Software Foundation.

Short Contents

1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 GnomeVFSResult . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 GnomeVFSURI. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 4 gnome-vfs-utils . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5 GnomeVFSFileInfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 6 Basic File Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 7 Truncating Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 8 Getting and Setting File Information . . . . . . . . . . . . . . . . . . . . 19 9 Basic Directory Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 10 GnomeVFSVolume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 11 GnomeVFSDrive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 12 GnomeVFSVolumeMonitor . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 13 MIME typing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 14 gnome-vfs-mime-utils . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 15 MIME Database . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 16 Undocumented . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Type Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 Function Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

1 Overview

(gnome vfs) wraps the GNOME Virtual Filesystem library for Guile. It is a part of Guile-GNOME.

The basic idea in this module is to map, as transparently as possible, GNOME-VFS files onto Scheme ports. Once you have a Scheme port, the native Scheme API is the preferred API. Other GNOME-VFS library procedures are defined that have no counterpart on the Scheme level, such as the MIME database procedures.

So, for example, to write to a file over SCP, you might do this:

     (define (with-output-to-port/dynamic make-port thunk)
       (let ((port #f))
         (dynamic-wind (lambda () (set! port (make-port)))
                       (lambda () (with-output-to-port port thunk))
                       (lambda () (close-port port) (set! port #f)))))
     (define (make-output-port uri-string exclusive?)
       (gnome-vfs-create
        (gnome-vfs-make-uri-from-input uri-string)
        'write exclusive? #o644))
     (with-output-to-port/dynamic
         (lambda ()
           (make-output-port "sftp://me@example.com/tmp/foo.txt" #t))
       (lambda ()
         (write "Hello world!")))

The dynamic-wind trickery is to ensure that the port is closed after execution leaves the thunk, and not left for GC to close in the future.

Exceptions raised during I/O are thrown to the gnome-vfs-error key, with the second argument being a symbol corresponding to a particular <gnome-vfs-result> value, such

as error-file-exists.

To enable integration with the GNOME keyring, for SSH keys and the like, you will need to call (gnome-authentication-manager-init), which is a procedure defined in the

(gnome gnome-ui) library.

This manual is admittedly a bit incomplete. Patches are accepted, of course, but the best thing to do would be eventually to wrap GIO, the new VFS layer that was pushed down into GLib.

See the documentation for (gnome gobject) for more information on Guile-GNOME.

2 GnomeVFSResult

Result of I/O operations, the equivalent of errno

2.1 Overview

2.2 Usage

[Function] gnome-vfs-result-to-string (res <gnome-vfs-result>)
   ⇒ (ret mchars)

Returns a string representing result, useful for debugging purposes, but probably not appropriate for passing to the user.

result a <gnome-vfs-result> to convert to a string.

ret a string representing result.

[Function] gnome-vfs-result-from-errno-code (errno code int)
   ⇒ (ret <gnome-vfs-result>)

Converts a system errno value to a <gnome-vfs-result>.

errno-code integer of the same type as the system "errno".

ret a <gnome-vfs-result> equivalent to errno-code.

[Function] gnome-vfs-result-from-errno ⇒ (ret <gnome-vfs-result>)

Converts the system errno to a <gnome-vfs-result>.

ret a <gnome-vfs-result> equivalent to the current system errno.

[Function] gnome-vfs-result-from-h-errno ⇒ (ret <gnome-vfs-result>)

Converts the system "h errno" to a <gnome-vfs-result> (h errno represents errors accessing and finding internet hosts)

ret a <gnome-vfs-result> equivalent to the current system "h errno".

[Function] gnome-vfs-result-from-h-errno-val (h errno code int)
   ⇒ (ret <gnome-vfs-result>)

Converts the error code h-errno-code into a <gnome-vfs-result>.

h-errno-code an integer representing the same error code as the system h errno.

ret The <gnome-vfs-result> equivalent to the h-errno-code.

3 GnomeVFSURI

A uniform resource identifier.

3.1 Overview

A <gnome-vfsuri> is a semi-textual representation of a uniform resource identifier. It contains the information about a resource location encoded as canononicalized text, but also holds extra information about the context in which the URI is used.

3.2 Usage

[Class] <gnome-vfsuri>

Opaque pointer. This class defines no direct slots.

[Function] gnome-vfs-uri-new (text uri mchars) ⇒ (ret <gnome-vfsuri>)

Create a new uri from text-uri. Unsupported and unsafe methods are not allowed and will result in ‘#f’ being returned. URL transforms are allowed. The a text uri must be an escaped URI string such as returned by gnome-vfs-get-

uri-from-local-path, gnome-vfs-make-uri-from-input, or gtk-file-chooser-
get-uri.

text-uri an escaped string representing a uri.

ret The new uri.

[Function] gnome-vfs-uri-resolve-relative (self <gnome-vfsuri>)

(relative reference mchars) ⇒ (ret <gnome-vfsuri>) Create a new uri from relative-reference, relative to base. The resolution algorithm in

some aspects follows RFC 2396 (http://www.ietf.org/rfc/rfc2396.txt), section

5.2, but is not identical due to some extra assumptions GnomeVFS makes about URIs. If relative-reference begins with a valid scheme identifier followed by ’:’, it is assumed to refer to an absolute URI, and a <gnome-vfsuri> is created from it using gnome-

vfs-uri-new.

Otherwise, depending on its precise syntax, it inherits some aspects of the parent URI, but the parents’ fragment and query components are ignored. If relative-reference begins with "//", it only inherits the base scheme, if it begins with ’/’ (i.e. is an absolute path reference), it inherits everything ecxept the base path. Otherwise, it replaces the part of base after the last ’/’.

This function should not be used by application authors unless they expect very distinct semantics. Instead, authors should use gnome-vfs-uri-append-file-name,

 gnome-vfs-uri-append-path, gnome-vfs-uri-append-string or gnome-vfs-uri-
 resolve-symbolic-link.

base base uri. relative-reference

a string representing a possibly relative uri reference.

ret A <gnome-vfsuri> referring to relative-reference, or ‘#f’ if relative-reference was malformed.

[Function] gnome-vfs-uri-resolve-symbolic-link (self <gnome-vfsuri>)

(relative reference mchars) ⇒ (ret <gnome-vfsuri>) Create a new uri from symbolic-link, relative to base. If symbolic-link begins with a ’/’, it replaces the path of base, otherwise it is appended after the last ’/’ character of base.

base base uri.

relative-reference ret A new <gnome-vfsuri> referring to symbolic-link.

Since 2.16

[Function] gnome-vfs-uri-append-string (self <gnome-vfsuri>)
   (uri fragment mchars) ⇒ (ret <gnome-vfsuri>)

Create a new uri obtained by appending uri-fragment to uri. This will take care of adding an appropriate directory separator between the end of uri and the start of uri-fragment if necessary.

uri  a <gnome-vfsuri>.

uri-fragment a piece of a uri (ie a fully escaped partial path).

ret The new uri obtained by combining uri and uri-fragment.

[Function] gnome-vfs-uri-append-path (self <gnome-vfsuri>) (path mchars)
   ⇒ (ret <gnome-vfsuri>)

Create a new uri obtained by appending path to uri. This will take care of adding an appropriate directory separator between the end of uri and the start of path if necessary as well as escaping path as necessary.

uri  a <gnome-vfsuri>.

path a non-escaped file path.

ret The new uri obtained by combining uri and path.

[Function] gnome-vfs-uri-append-file-name (self <gnome-vfsuri>)
   (filename mchars) ⇒ (ret <gnome-vfsuri>)

Create a new uri obtained by appending file-name to uri. This will take care of adding an appropriate directory separator between the end of uri and the start of file-name if necessary. file-name might, for instance, be the result of a call to g-dir-read-name.

uri  a <gnome-vfsuri>.

filename any "regular" file name (can include #, /, etc) in the file system encoding. This is not an escaped URI.

ret The new uri obtained by combining uri and path.

[Function] gnome-vfs-uri-to-string (self <gnome-vfsuri>)
   (hide options <gnome-vfsuri-hide-options>) ⇒ (ret mchars)

Translate uri into a printable string. The string will not contain the uri elements specified by hide-options. A file: URI on Win32 might look like file:///x:/foo/bar.txt. Note that the part after file:// is not a legal file name, you need to remove the / in front of the drive letter. This function does that automatically if hide-options specifies that the toplevel method, user name, password, host name and host port should be hidden. On the other hand, a file: URI for a UNC path looks like file:////server/share/foo/bar.txt, and in that case the part after file:// is the correct file name.

uri  a <gnome-vfsuri>.

hide-options bitmask specifying what uri elements (e.g. password, user name etc.) should not be represented in the returned string.

ret a malloc’d printable string representing uri.

[Function] gnome-vfs-uri-dup (self <gnome-vfsuri>) ⇒ (ret <gnome-vfsuri>)

Duplicate uri.

uri  a <gnome-vfsuri>.

ret a pointer to a new uri that is exactly the same as uri.

[Function] gnome-vfs-uri-is-local (self <gnome-vfsuri>) ⇒ (ret bool)

Check if uri is a local URI. Note that the return value of this function entirely depends on the <gnome-vfs-method> associated with the URI. It is up to the method author to distinguish between remote URIs add URIs referring to entities on the local computer. Warning, this can be slow, as it does i/o to detect things like NFS mounts.

uri  a <gnome-vfsuri>.

ret ‘#t’ if uri refers to a local entity, ‘#f’ otherwise.

[Function] gnome-vfs-uri-has-parent (self <gnome-vfsuri>) ⇒ (ret bool)

Check if uri has a parent or not.

uri  a <gnome-vfsuri>.

ret ‘#t’ if uri has a parent, ‘#f’ otherwise.

[Function] gnome-vfs-uri-get-parent (self <gnome-vfsuri>)
   ⇒ (ret <gnome-vfsuri>)

Retrieve uri’s parent uri.

uri  a <gnome-vfsuri>.

ret a pointer to uri’s parent uri.

[Function] gnome-vfs-uri-get-host-name (self <gnome-vfsuri>)
   ⇒ (ret mchars)

Retrieve the host name for uri.

uri  a <gnome-vfsuri>.

ret a string representing the host name.

[Function] gnome-vfs-uri-get-scheme (self <gnome-vfsuri>) ⇒ (ret mchars)

Retrieve the scheme used for uri.

uri  a <gnome-vfsuri>.

ret a string representing the scheme.

[Function] gnome-vfs-uri-get-host-port (self <gnome-vfsuri>)
   ⇒ (ret unsigned-int)

Retrieve the host port number in uri.

uri  a <gnome-vfsuri>.

ret The host port number used by uri. If the value is zero, the default port

value for the specified toplevel access method is used.

[Function] gnome-vfs-uri-get-user-name (self <gnome-vfsuri>)
   ⇒ (ret mchars)

Retrieve the user name in uri.

uri  a <gnome-vfsuri>.

ret a string representing the user name in uri.

[Function] gnome-vfs-uri-get-password (self <gnome-vfsuri>)
   ⇒ (ret mchars)

Retrieve the password for uri.

uri  a <gnome-vfsuri>.

ret The password for uri.

[Function] gnome-vfs-uri-set-host-name (self <gnome-vfsuri>)

(host name mchars) Set host-name as the host name accessed by uri.

uri  a <gnome-vfsuri>.

host-name a string representing a host name.

[Function] gnome-vfs-uri-set-host-port (self <gnome-vfsuri>)
   (host port unsigned-int)

Set the host port number in uri. If host-port is zero, the default port for uri’s toplevel access method is used.

uri  a <gnome-vfsuri>.

host-port a TCP/IP port number.

[Function] gnome-vfs-uri-set-user-name (self <gnome-vfsuri>)

(user name mchars) Set user-name as the user name for uri.

uri  a <gnome-vfsuri>.

user-name a string representing a user name on the host accessed by uri.

[Function] gnome-vfs-uri-set-password (self <gnome-vfsuri>)

(password mchars) Set password as the password for uri.

uri  a <gnome-vfsuri>.

password a password string.

[Function] gnome-vfs-uri-equal (self <gnome-vfsuri>) (b <gnome-vfsuri>)

⇒ (ret bool) Compare a and b.

FIXME: This comparison should take into account the possiblity that unreserved characters may be escaped. ...or perhaps gnome-vfs-uri-new should unescape unreserved characters?

a  a <gnome-vfsuri>.
b  a <gnome-vfsuri>.

ret ‘#t’ if a and b are equal, ‘#f’ otherwise.

[Function] gnome-vfs-uri-is-parent (self <gnome-vfsuri>)

(possible child <gnome-vfsuri>) (recursive bool) ⇒ (ret bool) Check if possible-child is contained by possible-parent. If recursive is ‘#f’, just try the immediate parent directory, else search up through the hierarchy.

possible-parent

          a <gnome-vfsuri>.

possible-child

          a <gnome-vfsuri>.

recursive a flag to turn recursive check on.

ret ‘#t’ if possible-child is contained in possible-parent.

[Function] gnome-vfs-uri-get-path (self <gnome-vfsuri>) ⇒ (ret mchars)

Retrieve full path name for uri.

uri  a <gnome-vfsuri>.

ret a pointer to the full path name in uri. Notice that the pointer points to the path name stored in uri, so the path name returned must not be modified nor freed.

[Function] gnome-vfs-uri-extract-dirname (self <gnome-vfsuri>)
   ⇒ (ret mchars)

Extract the name of the directory in which the file pointed to by uri is stored as a newly allocated string. The string will end with a ‘GNOME_VFS_URI_PATH_CHR’.

uri  a <gnome-vfsuri>.

ret a pointer to the newly allocated string representing the parent directory.

[Function] gnome-vfs-uri-extract-short-name (self <gnome-vfsuri>)
   ⇒ (ret mchars)

Retrieve base file name for uri, ignoring any trailing path separators. This matches the XPG definition of basename, but not g basename. This is often useful when you want the name of something that’s pointed to by a uri, and don’t care whether the uri has a directory or file form. If uri points to the root of a domain, returns the host name. If there’s no host name, returns ‘GNOME_VFS_URI_PATH_STR’.

See also: gnome-vfs-uri-extract-short-path-name.
uri  a <gnome-vfsuri>.

ret a pointer to the newly allocated string representing the unescaped short form of the name.

[Function] gnome-vfs-uri-list-parse (uri list mchars) ⇒ (ret glist-of)

Extracts a list of <gnome-vfsuri> objects from a standard text/uri-list, such as one you would get on a drop operation. Use gnome-vfs-uri-list-free when you are done with the list.

uri-list string consists of <gnome-vfsur-is> and/or paths seperated by newline character.

ret  a <g-list> of <gnome-vfsur-is>.

4 gnome-vfs-utils

various utilities functions to manipulate uris

4.1 Overview

4.2 Usage

[Function] gnome-vfs-format-uri-for-display (uri mchars) ⇒ (ret mchars)

Filter, modify, unescape and change uri to make it appropriate for display to users. The conversion is done such that the roundtrip to UTF-8 is reversible.

Rules: file: uri without fragments should appear as local paths. file: uri with fragments should appear as file:uri. All other uri appear as expected.

uri uri to be displayed.

ret a string which represents uri and can be displayed.

Since 2.2

[Function] gnome-vfs-url-show (url mchars) ⇒ (ret <gnome-vfs-result>)

Launches the default application or component associated with the given url.

url url to be shown.

ret  ‘GNOME_VFS_OK’ if the default action was launched, ‘GNOME_VFS_ERROR_BAD_PARAMETERS’

for an invalid or non-existant url, ‘GNOME_VFS_ERROR_NO_DEFAULT’ if no default action is associated with the url. Also error codes from

          gnome-vfs-mime-action-launch or gnome-vfs-mime-action-launch-
          with-env.

Since 2.4

[Function] gnome-vfs-escape-string (string mchars) ⇒ (ret mchars)

Escapes string, replacing any and all special characters with equivalent escape sequences.

string string to be escaped.

ret a newly allocated string equivalent to string but with all special characters escaped.

[Function] gnome-vfs-escape-path-string (path mchars) ⇒ (ret mchars)

Escapes path, replacing only special characters that would not be found in paths (so ’/’, ’&’, and ’=’ will not be escaped by this function).

path string to be escaped.

ret a newly allocated string equivalent to path but with non-path characters escaped.

[Function] gnome-vfs-escape-slashes (string mchars) ⇒ (ret mchars)

Escapes only ’/’ and ’%’ characters in string, replacing them with their escape sequence equivalents.

string string to be escaped.

ret a newly allocated string equivalent to string, but with no unescaped ’/’ or ’%’ characters.

[Function] gnome-vfs-escape-set (string mchars) (match set mchars)
   ⇒ (ret mchars)

Escapes all characters in string which are listed in match-set.

string string to be escaped.

match-set a string containing all characters to be escaped in string.

ret a newly allocated string equivalent to string but with characters in matchstring escaped.

[Function] gnome-vfs-unescape-string (escaped string mchars)

(illegal characters mchars) ⇒ (ret mchars) Decodes escaped characters (i.e. PERCENTxx sequences) in escaped-string. Characters are encoded in PERCENTxy form, where xy is the ASCII hex code for character 16x+y.

escaped-string an escaped uri, path, or other string.

illegal-characters a string containing a sequence of characters considered "illegal" to be escaped, ’\0’ is automatically in this list.

ret a newly allocated string with the unescaped equivalents, or ‘#f’ if escapedstring contained an escaped encoding of one of the characters in illegalcharacters.

[Function] gnome-vfs-make-uri-canonical (uri mchars) ⇒ (ret mchars)

Standardizes the format of the uri, so that it can be used later in other functions that expect a canonical uri.

uri an absolute or relative stringified uri. It might have scheme.

ret a newly allocated string that contains the canonical representation of uri.

Since 2.2

[Function] gnome-vfs-make-path-name-canonical (path mchars)
   ⇒ (ret mchars)

Calls -gnome-vfs-canonicalize-pathname, allocating storage for the result and providing for a cleaner memory management.

path a file path, relative or absolute.

ret a canonical version of path.

[Function] gnome-vfs-make-uri-from-input (location mchars) ⇒ (ret mchars)

Takes a user input path/uri and makes a valid uri out of it. This function is the reverse of gnome-vfs-format-uri-for-display but it also handles the fact that the user could have typed arbitrary UTF-8 in the entry showing the string.

location a possibly mangled "uri", in UTF-8.

ret a newly allocated uri.

Since 2.2

[Function] gnome-vfs-make-uri-from-shell-arg (uri mchars) ⇒ (ret mchars)
Similar to gnome-vfs-make-uri-from-input, except that:

1) guesses relative paths instead of http domains. 2) doesn’t bother stripping leading/trailing white space. 3) doesn’t bother with ~ expansion–that’s done by the shell.

uri path to make the uri from.

ret a newly allocated string representing uri.

Since 2.2

[Function] gnome-vfs-expand-initial-tilde (path mchars) ⇒ (ret mchars)

If path starts with a ~, representing the user’s home directory, expand it to the actual path location.

path a local file path which may start with a ’~’.

ret a newly allocated string with the initial tilde (if there was one) converted to an actual path.

[Function] gnome-vfs-get-local-path-from-uri (uri mchars) ⇒ (ret mchars)

Create a local path for a file:/// uri. Do not use with uris of other methods.

uri uri to convert to a local path.

ret a newly allocated string containing the local path. ‘#f’ is returned on error or if the uri isn’t a file: uri without a fragment identifier (or chained uri).

[Function] gnome-vfs-get-uri-from-local-path (local full path mchars)
   ⇒ (ret mchars)

Returns a file:/// URI for the local path local-full-path, such as a path provided by gtk-file-chooser-get-filename. The resulting URI may be provided, for instance,

to gnome-vfs-uri-new.

On Windows local-full-path should be in the UTF-8 encoding, and can start with a drive letter, but doesn’t have to.

local-full-path a full local filesystem path (i.e. not relative).

ret a newly allocated string containing the uri corresponding to local-full-path (‘#f’ for some bad errors).

[Function] gnome-vfs-icon-path-from-filename (filename mchars)
   ⇒ (ret mchars)

filename path to a file. Could be relative or absolute path. ret Returns the icon path for the filename. Example: gnome vfs icon path from filename

("nautilus/nautilus-desktop.png") will return a string forming the full path of the file nautilus-desktop.png ie $PREFIX/share/pixmaps/nautilus/nautilus-desktop.png.

[Function] gnome-vfs-is-primary-thread ⇒ (ret bool)

Check if the current thread is the thread with the main glib event loop. ret ‘#t’ if the current thread is the thread with the main glib event loop.

[Function] gnome-vfs-get-uri-scheme (uri mchars) ⇒ (ret mchars)

Retrieve the scheme used in uri. uri a stringified uri. ret a newly allocated string containing the scheme, ‘#f’ if uri doesn’t contain

a scheme.

Since 2.2

[Function] gnome-vfs-uris-match (uri 1 mchars) (uri 2 mchars) ⇒ (ret bool)

Compare two uris. uri-1 stringified uri to compare with uri-2. uri-2 stringified uri to compare with uri-1. ret ‘#t’ if they are the same, ‘#f’ otherwise. Since 2.2 Chapter 5: GnomeVFSFileInfo 13

5 GnomeVFSFileInfo

stores information about files, GnomeVFS equivalent of stat

5.1 Overview 5.2 Usage

[Class] <gnome-vfs-file-info>

Opaque pointer.

This class defines no direct slots.

6 Basic File Operations

Essential VFS operations. This includes creating, moving and deleting files, opening and closing file handles.

6.1 Overview

GnomeVFS file operations are, for the most part, patterned after their POSIX equivalents, with the systematic difference that they accept URIs rather than paths on the local filesystem. This makes them easy to learn as if you are already familiar with basic commands such as open, seek, write, etc you will feel right at home with GnomeVFS after learning a little about URIs.

6.2 Usage

[Function] gnome-vfs-create (text uri mchars)
   (open mode <gnome-vfs-open-mode>) (exclusive bool)
   (perm unsigned-int) ⇒ (ret scm)

Create text-uri according to mode open-mode. On return, handle will then contain a pointer to a handle for the open file.

handle pointer to a pointer to a <gnome-vfs-handle> object.

text-uri string representing the uri to create.

open-mode mode to leave the file opened in after creation (or ‘GNOME_VFS_OPEN_MODE_NONE’ to leave the file closed after creation).

exclusive whether the file should be created in "exclusive" mode. i.e. if this flag is nonzero, operation will fail if a file with the same name already exists.

perm bitmap representing the permissions for the newly created file (Unix style).

ret an integer representing the result of the operation.

[Function] gnome-vfs-create-uri (uri <gnome-vfsuri>)
   (open mode <gnome-vfs-open-mode>) (exclusive bool)
   (perm unsigned-int) ⇒ (ret scm)

Create uri according to mode open-mode. On return, handle will then contain a pointer to a handle for the open file.

handle pointer to a pointer to a <gnome-vfs-handle> object.

uri uri for the file to create.

open-mode open mode.

exclusive whether the file should be created in "exclusive" mode. i.e. if this flag is nonzero, operation will fail if a file with the same name already exists. perm bitmap representing the permissions for the newly created file (Unix

style).

ret an integer representing the result of the operation.

[Function] gnome-vfs-open (text uri mchars)
   (open mode <gnome-vfs-open-mode>) ⇒ (ret scm)

Open text-uri according to mode open-mode. On return, handle will then contain a pointer to a handle for the open file.

handle pointer to a pointer to a <gnome-vfs-handle> object.

text-uri string representing the uri to open.

open-mode open mode.

ret an integer representing the result of the operation.

[Function] gnome-vfs-open-uri (uri <gnome-vfsuri>)
   (open mode <gnome-vfs-open-mode>) ⇒ (ret scm)

Open uri according to mode open-mode. On return, handle will then contain a pointer to a handle for the open file.

handle pointer to a pointer to a <gnome-vfs-handle> object.

uri uri to open.

open-mode open mode.

ret an integer representing the result of the operation.

[Function] gnome-vfs-unlink (text uri mchars) ⇒ (ret <gnome-vfs-result>)

Unlink text-uri (i.e. delete the file).

text-uri uri of the file to be unlinked.

ret an integer representing the result of the operation.

[Function] gnome-vfs-unlink-from-uri (uri <gnome-vfsuri>)
   ⇒ (ret <gnome-vfs-result>)

Unlink uri (i.e. delete the file).

uri uri of the file to be unlinked.

ret an integer representing the result of the operation.

[Function] gnome-vfs-move-uri (old uri <gnome-vfsuri>)

(new uri <gnome-vfsuri>) (force replace bool)

   ⇒ (ret <gnome-vfs-result>)

Move a file from uri old-uri to new-uri. This will only work if old-uri and new-uri are on the same file system. Otherwise, it is necessary to use the more general gnome-

vfs-xfer-uri function.

old-uri source uri. new-uri destination uri.

force-replace if ‘#t’, move old-uri to new-uri even if there is already a file at new-uri. If there is a file, it will be discarded.

ret an integer representing the result of the operation.

[Function] gnome-vfs-move (old text uri mchars) (new text uri mchars)

   (force replace bool) ⇒ (ret <gnome-vfs-result>)

Move a file from old-text-uri to new-text-uri. This will only work if old-text-uri and new-text-uri are on the same file system. Otherwise, it is necessary to use the more

general gnome-vfs-xfer-uri function.

old-text-uri string representing the source file location.

new-text-uri string representing the destination file location.

force-replace if ‘#t’, perform the operation even if it unlinks an existing file at newtext-uri.

ret an integer representing the result of the operation.

[Function] gnome-vfs-check-same-fs-uris (source uri <gnome-vfsuri>)
   (target uri <gnome-vfsuri>) ⇒ (ret <gnome-vfs-result>)

(same fs return bool) Check if source-uri and target-uri are on the same file system.

source-uri a uri.

target-uri another uri.

same-fs-return pointer to a boolean variable which will be set to ‘#t’ on return if sourceuri and target-uri are on the same file system.

ret an integer representing the result of the operation.

[Function] gnome-vfs-check-same-fs (source mchars) (target mchars)
   ⇒ (ret <gnome-vfs-result>) (same fs return bool)

Check if source and target are on the same file system.

source path to a file.

target path to another file.

same-fs-return pointer to a boolean variable which will be set to ‘#t’ on return if source and target are on the same file system.

ret an integer representing the result of the operation.

[Function] gnome-vfs-uri-exists (self <gnome-vfsuri>) ⇒ (ret bool)

Check if the uri points to an existing entity.

uri a uri.

ret ‘#t’ if uri exists.

[Function] gnome-vfs-create-symbolic-link (uri <gnome-vfsuri>)
   (target reference mchars) ⇒ (ret <gnome-vfs-result>)

Creates a symbolic link, or eventually, a uri link (as necessary) at uri pointing to target-reference.

uri uri to create a link at.

target-reference uri "reference" to point the link to (uri or relative path).

ret an integer representing the result of the operation.

7 Truncating Files

Force files to a particular length

7.1 Overview

Truncation of files is used to force them to a particular length. If a file longer than specified, the trailing bytes are discarded, if it is shorter than specified it is padded with zeros.

7.2 Usage

[Function] gnome-vfs-truncate (text uri mchars) (length unsigned-int64)
   ⇒ (ret <gnome-vfs-result>)

Truncate the file at text-uri to length bytes. text-uri string representing the file to be truncated. length length of the new file at text-uri. ret an integer representing the result of the operation.

[Function] gnome-vfs-truncate-uri (uri <gnome-vfsuri>)
   (length unsigned-int64) ⇒ (ret <gnome-vfs-result>)

Truncate the file at uri to be only length bytes. Data past length bytes will be discarded. uri uri of the file to be truncated. length length of the new file at uri. ret an integer representing the result of the operation.

8 Getting and Setting File Information

Convenient high-level abstraction for obtaining and setting file information, including ACLs.

8.1 Overview

Applications can use the gnome-vfs-get-file-info family of operations to retrieve file information, as this operation can be quite costly in terms of time (specilly when sniffing the MIME type) applications can specify which information need at any time, reducing the performance impact.

All of these operations use a <gnome-vfs-file-info> data structure that holds the file information, there are several methods that can be used to manipulate this information. See <gnome-vfs-file-info> for more information.

8.2 Usage

[Function] gnome-vfs-get-file-info (text uri mchars)
   (info <gnome-vfs-file-info>)
   (options <gnome-vfs-file-info-options>)
   ⇒ (ret <gnome-vfs-result>)

Retrieve information about text-uri. The information will be stored in info.

text-uri uri of the file for which information will be retrieved.

info pointer to a <gnome-vfs-file-info> object that will hold the information for the file on return.

options options for retrieving file information.

ret an integer representing the result of the operation.

[Function] gnome-vfs-set-file-info (text uri mchars)
   (info <gnome-vfs-file-info>)
   (mask <gnome-vfs-set-file-info-mask>)
   ⇒ (ret <gnome-vfs-result>)

Set file information for uri; only the information for which the corresponding bit in mask is set is actually modified.

 info’s ‘valid_fields’ is not required to contain the <gnome-vfs-file-info-fields>

corresponding to the specified <gnome-vfs-set-file-info-mask> fields of mask. It is assumed that the info fields referenced by mask are valid.

text-uri string representing the file location.

info information that must be set for the file.

mask bit mask representing which fields of info need to be set.

ret an integer representing the result of the operation.

9 Basic Directory Operations

Creating and removing directories.

9.1 Overview

9.2 Usage

[Function] gnome-vfs-make-directory (text uri mchars) (perm unsigned-int)
   ⇒ (ret <gnome-vfs-result>)

Create text-uri as a directory. text-uri uri of the directory to be created. perm Unix-style permissions for the newly created directory ret an integer representing the result of the operation.

[Function] gnome-vfs-make-directory-for-uri (uri <gnome-vfsuri>)
   (perm unsigned-int) ⇒ (ret <gnome-vfs-result>)

Create a directory at uri. Only succeeds if a file or directory does not already exist at uri. uri uri of the directory to be created. perm Unix-style permissions for the newly created directory. ret an integer representing the result of the operation.

[Function] gnome-vfs-remove-directory (text uri mchars)
   ⇒ (ret <gnome-vfs-result>)

Remove text-uri. text-uri must be an empty directory. text-uri path of the directory to be removed. ret an integer representing the result of the operation.

[Function] gnome-vfs-remove-directory-from-uri (uri <gnome-vfsuri>)
   ⇒ (ret <gnome-vfs-result>)

Remove uri. uri must be an empty directory. uri uri of the directory to be removed. ret an integer representing the result of the operation.

10 GnomeVFSVolume

Abstraction for a mounted file system or a network location.

10.1 Overview

10.2 Usage

[Class] <gnome-vfs-volume>

Derives from <gobject>. This class defines no direct slots.

[Function] gnome-vfs-volume-compare (self <gnome-vfs-volume>)
   (b <gnome-vfs-volume>) ⇒ (ret int)

[Method] compare

Compares two <gnome-vfs-volume> objects a and b. Two <gnome-vfs-volume>

objects referring to different volumes are guaranteed to not return 0 when comparing them, if they refer to the same volume 0 is returned. The resulting <gint> should be used to determine the order in which a and b are displayed in graphical user interfces. The comparison algorithm first of all peeks the device type of a and b, they will be sorted in the following order:

• • • • • •

Magnetic and opto-magnetic volumes (ZIP, floppy) Optical volumes (CD, DVD) External volumes (USB sticks, music players) Mounted hard disks Network mounts Other volumes Afterwards, the display name of a and b is compared using a locale-sensitive sorting algorithm, which involves g-utf8-collate-key. If two volumes have the same display name, their unique ID is compared which can

be queried using gnome-vfs-volume-get-id.
a  a <gnome-vfs-volume>.
b  a <gnome-vfs-volume>.

ret 0 if the volumes refer to the same gnome-vfs-volume, a negative value if a should be displayed before b, or a positive value if a should be displayed after b.

Since 2.6

[Function] gnome-vfs-volume-get-activation-uri (self <gnome-vfs-volume>)
   ⇒ (ret mchars)
[Method] get-activation-uri

Returns the activation URI of a <gnome-vfs-volume>. The returned URI usually refers to a valid location. You can check the validity of the location by calling gnome-vfs-uri-new with the URI, and checking whether the return value is not ‘#f’.

volume  a <gnome-vfs-volume>.

ret a newly allocated string for activation uri of volume.

Since 2.6

[Function] gnome-vfs-volume-get-device-path (self <gnome-vfs-volume>)
   ⇒ (ret mchars)
[Method] get-device-path

Returns the device path of a <gnome-vfs-volume>. For HAL volumes, this returns the value of the volume’s "block.device" key. For UNIX mounts, it returns the ‘mntent’’s ‘mnt_fsname’ entry. Otherwise, it returns ‘#f’.

volume  a <gnome-vfs-volume>.

ret a newly allocated string for device path of volume.

Since 2.6

[Function] gnome-vfs-volume-get-device-type (self <gnome-vfs-volume>)
   ⇒ (ret <gnome-vfs-device-type>)
[Method] get-device-type
Returns the <gnome-vfs-device-type> of the volume.
volume  a <gnome-vfs-volume>.

ret the device type for volume.

Since 2.6

[Function] gnome-vfs-volume-get-display-name (self <gnome-vfs-volume>)
   ⇒ (ret mchars)
[Method] get-display-name

Returns the display name of the volume.

volume  a <gnome-vfs-volume>.

ret a newly allocated string for display name of volume.

Since 2.6

[Function] gnome-vfs-volume-get-drive (self <gnome-vfs-volume>)
   ⇒ (ret <gnome-vfs-drive>)
[Method] get-drive
volume  a <gnome-vfs-volume>.

ret the drive for the volume.

Since 2.6

[Function] gnome-vfs-volume-get-hal-udi (self <gnome-vfs-volume>)
   ⇒ (ret mchars)
[Method] get-hal-udi

Returns the HAL UDI of a <gnome-vfs-volume>. For HAL volumes, this matches the value of the "info.udi" key, for other volumes it is ‘#f’.

volume  a <gnome-vfs-volume>.

ret a newly allocated string for unique device id of volume, or ‘#f’.

Since 2.6

[Function] gnome-vfs-volume-get-icon (self <gnome-vfs-volume>)
   ⇒ (ret mchars)
[Method] get-icon
volume  a <gnome-vfs-volume>.

ret a newly allocated string for the icon filename of volume.

Since 2.6

[Function] gnome-vfs-volume-get-id (self <gnome-vfs-volume>)
   ⇒ (ret unsigned-long)

[Method] get-id

Returns the ID of the volume, Two <gnome-vfs-volumes> are guaranteed to refer to the same volume if they have the same ID.

volume  a <gnome-vfs-volume>.

ret the id for the volume.

Since 2.6

[Function] gnome-vfs-volume-get-volume-type (self <gnome-vfs-volume>)
   ⇒ (ret <gnome-vfs-volume-type>)
[Method] get-volume-type
Returns the <gnome-vfs-volume-type> of the volume.
volume  a <gnome-vfs-volume>.

ret the volume type for volume.

Since 2.6

[Function] gnome-vfs-volume-handles-trash (self <gnome-vfs-volume>)

⇒ (ret bool)

[Method] handles-trash

Returns whether the file system on a volume supports trashing of files. If the volume has an AutoFS file system (i.e. gnome-vfs-volume-get-device-type returns <gnome-vfs-device-type-autofs>), or if the volume is mounted read-only (gnome-vfs-volume-is-read-only returns ‘#t’), it is assumed to not support trashing of files. Otherwise, if the volume provides file system information, it is determined whether the file system supports trashing of files. See gnome-vfs-volume-get-filesystem-type for details which volumes provide file system information, and which file systems currently support a trash.

volume  a <gnome-vfs-volume>.

ret ‘#t’ if volume handles trash, ‘#f’ otherwise.

Since 2.6

[Function] gnome-vfs-volume-is-mounted (self <gnome-vfs-volume>)

⇒ (ret bool)

[Method] is-mounted

Returns whether the file system on a volume is currently mounted. For HAL volumes, this reflects the value of the "volume.is mounted" key, for traditional UNIX mounts and connected servers, ‘#t’ is returned, because their existence implies that they are mounted.

volume  a <gnome-vfs-volume>.

ret ‘#t’ if the volume is mounted, ‘#f’ otherwise.

Since 2.6

[Function] gnome-vfs-volume-is-read-only (self <gnome-vfs-volume>)

⇒ (ret bool)

[Method] is-read-only

Returns whether the file system on a volume is read-only. For HAL volumes, the "volume.is mounted read only" key is authoritative, for traditional UNIX mounts it returns ‘#t’ if the mount was done with the "ro" option. For servers, ‘#f’ is returned.

volume  a <gnome-vfs-volume>.

ret ‘#t’ if the volume is read-only to the user, ‘#f’ otherwise.

Since 2.6

[Function] gnome-vfs-volume-is-user-visible (self <gnome-vfs-volume>)

⇒ (ret bool)

[Method] is-user-visible

Returns whether the volume is visible to the user. This should be used by applications to determine whether it is included in user interfaces listing available volumes.

volume  a <gnome-vfs-volume>.

ret ‘#t’ if volume is visible to the user, ‘#f’ otherwise.

Since 2.6

[Function] gnome-vfs-connect-to-server (uri mchars) (display name mchars)
   (icon mchars)

This function adds a server connection to the specified uri, which is displayed in user interfaces with the specified display-name and icon. If this function is invoked successfully, the created server shows up in the list of mounted volumes of the <gnome-vfs-volume-monitor>, which can be queried using

gnome-vfs-volume-monitor-get-mounted-volumes.

This function does not have a return value. Hence, you can’t easily detect whether the specified server was successfully created. The actual creation and consumption of the new server through the <gnome-vfs-volume-monitor> is done asynchronously. uri, display-name, and icon can be freely chosen, but should be meaningful: uri should refer to a valid location. You can check the validity of the location by calling gnome-vfs-uri-new with uri, and checking whether the return value is not ‘#f’. The display-name should be queried from the user, and an empty string should not be considered valid. icon typically references an icon from the icon theme. Some implemen-

tations currently use ‘gnome-fs-smb’, ‘gnome-fs-ssh’, ‘gnome-fs-ftp’

and ‘gnome-fs-share’, depending on the type of the server referenced by uri. The icon naming conventions might change in the future, though. Obeying the freedesktop.org Icon Naming Specification

(http://standards.freedesktop.org/icon-naming-spec/icon-naming-spec-latest.

html) is suggested.

uri The string representation of the server to connect to.

display-name The display name that is used to identify the server connection.

icon The icon that is used to identify the server connection.

Since 2.6

11 GnomeVFSDrive

Container for GnomeVFSVolume (floppy drive, CD reader, ...)

11.1 Overview

11.2 Usage

[Class] <gnome-vfs-drive>

Derives from <gobject>. This class defines no direct slots.

[Signal on <gnome-vfs-drive>] volume-mounted (arg0 <gnome-vfs-volume>)

This signal is emitted after the <gnome-vfs-volume>volume has been mounted. When the volume is mounted, it is added to the drive’s list of mounted volumes,

which can be queried using gnome-vfs-drive-get-mounted-volumes.

It is also added to the list of the <gnome-vfs-volume-monitor>’s list of mounted

volumes, which can be queried using gnome-vfs-volume-monitor-get-mounted-
volumes.
[Signal on <gnome-vfs-drive>] volume-pre-unmount
   (arg0 <gnome-vfs-volume>)

This signal is emitted when the <gnome-vfs-volume>volume, which has been present in the <gnome-vfs-drive>drive, is about to be unmounted. When the volume is unmounted, it is removed from the drive’s list of mounted vol-

umes, which can be queried using gnome-vfs-drive-get-mounted-volumes.

It is also removed from the <gnome-vfs-volume-monitor>’s list of mounted volumes,

which can be queried using gnome-vfs-volume-monitor-get-mounted-volumes.

When a client application receives this signal, it must free all resources associated with the volume, for instance cancel all pending file operations on the volume, and cancel all pending file monitors using gnome-vfs-monitor-cancel.

[Signal on <gnome-vfs-drive>] volume-unmounted (arg0 <gnome-vfs-volume>)

This signal is emitted after the <gnome-vfs-volume>volume, which had been present in the <gnome-vfs-drive>drive, has been unmounted. When the volume is unmounted, it is removed from the drive’s list of mounted vol-

umes, which can be queried using gnome-vfs-drive-get-mounted-volumes.

It is also removed from the <gnome-vfs-volume-monitor>’s list of mounted volumes,

which can be queried using gnome-vfs-volume-monitor-get-mounted-volumes.
[Function] gnome-vfs-drive-compare (self <gnome-vfs-drive>)
   (b <gnome-vfs-drive>) ⇒ (ret int)

[Method] compare

Compares two <gnome-vfs-drive> objects a and b. Two <gnome-vfs-drive> ob-

jects referring to different drives are guaranteed to not return 0 when comparing them, if they refer to the same drive 0 is returned.

The resulting <gint> should be used to determine the order in which a and b are displayed in graphical user interfces. The comparison algorithm first of all peeks the device type of a and b, they will be sorted in the following order:

• • • • •

Magnetic and opto-magnetic drives (ZIP, floppy) Optical drives (CD, DVD) External drives (USB sticks, music players) Mounted hard disks< Other drives< Afterwards, the display name of a and b is compared using a locale-sensitive sorting algorithm, which involves g-utf8-collate-key. If two drives have the same display name, their unique ID is compared which can be

queried using gnome-vfs-drive-get-id.
a  a <gnome-vfs-drive>.
b  a <gnome-vfs-drive>.

ret 0 if the drives refer to the same gnome-vfs-drive, a negative value if a should be displayed before b, or a positive value if a should be displayed after b.

Since 2.6

[Function] gnome-vfs-drive-get-activation-uri (self <gnome-vfs-drive>)
   ⇒ (ret mchars)
[Method] get-activation-uri

Returns the activation URI of a <gnome-vfs-drive>. The returned URI usually refers to a valid location. You can check the validity of the location by calling gnome-vfs-uri-new with the URI, and checking whether the return value is not ‘#f’.

drive  a <gnome-vfs-drive>.

ret a newly allocated string for the activation uri of the <drive>.

Since 2.6

[Function] gnome-vfs-drive-get-device-path (self <gnome-vfs-drive>)
   ⇒ (ret mchars)
[Method] get-device-path

Returns the device path of a <gnome-vfs-drive>.

For HAL drives, this returns the value of the drives’s "block.device" key. For UNIX mounts, it returns the ‘mntent’’s ‘mnt_fsname’ entry. Otherwise, it returns ‘#f’.

drive  a <gnome-vfs-drive>.

ret a newly allocated string for the device path of the <drive>.

Since 2.6

[Function] gnome-vfs-drive-get-device-type (self <gnome-vfs-drive>)
   ⇒ (ret <gnome-vfs-device-type>)
[Method] get-device-type
drive  a <gnome-vfs-drive>.
ret  device type, a <gnome-vfs-device-type> value.

Since 2.6

[Function] gnome-vfs-drive-get-display-name (self <gnome-vfs-drive>)
   ⇒ (ret mchars)
[Method] get-display-name
drive  a <gnome-vfs-drive>.

ret a newly allocated string for the display name of the drive.

Since 2.6

[Function] gnome-vfs-drive-get-hal-udi (self <gnome-vfs-drive>)
   ⇒ (ret mchars)
[Method] get-hal-udi

Returns the HAL UDI of a <gnome-vfs-drive>. For HAL drives, this matches the value of the "info.udi" key, for other drives it is ‘#f’.

drive  a <gnome-vfs-drive>.

ret a newly allocated string for the unique device id of the drive, or ‘#f’.

Since 2.6

[Function] gnome-vfs-drive-get-icon (self <gnome-vfs-drive>)
   ⇒ (ret mchars)
[Method] get-icon
drive  a <gnome-vfs-drive>.

ret a newly allocated string for the icon filename of the drive.

Since 2.6

[Function] gnome-vfs-drive-get-id (self <gnome-vfs-drive>)
   ⇒ (ret unsigned-long)

[Method] get-id

drive  a <gnome-vfs-drive>.

ret drive id, a <gulong> value. Since 2.6

[Function] gnome-vfs-drive-get-mounted-volumes (self <gnome-vfs-drive>)
   ⇒ (ret glist-of)
[Method] get-mounted-volumes
drive  a <gnome-vfs-drive>.

ret list of mounted volumes for the drive. Since 2.8

[Function] gnome-vfs-drive-is-connected (self <gnome-vfs-drive>)

⇒ (ret bool)

[Method] is-connected
drive  a <gnome-vfs-drive>.

ret ‘#t’ if the drive is connected, ‘#f’ otherwise. Since 2.6

[Function] gnome-vfs-drive-is-mounted (self <gnome-vfs-drive>)

⇒ (ret bool)

[Method] is-mounted
drive  a <gnome-vfs-drive>.

ret ‘#t’ if the drive is mounted, ‘#f’ otherwise. Since 2.6

[Function] gnome-vfs-drive-is-user-visible (self <gnome-vfs-drive>)

⇒ (ret bool)

[Method] is-user-visible

Returns whether the drive is visible to the user. This should be used by applications to determine whether it is included in user interfaces listing available drives.

drive  a <gnome-vfs-drive>.

ret ‘#t’ if the drive is visible to the user, ‘#f’ otherwise. Since 2.6

12 GnomeVFSVolumeMonitor

Monitors volume mounts and unmounts

12.1 Overview

12.2 Usage

[Class] <gnome-vfs-volume-monitor>

Derives from <gobject>. This class defines no direct slots.

[Signal on <gnome-vfs-volume-monitor>] volume-mounted
   (arg0 <gnome-vfs-volume>)

This signal is emitted after the <gnome-vfs-volume>volume has been mounted. When the volume is mounted, it is present in the volume-monitor’s list of mounted

volumes, which can be queried using gnome-vfs-volume-monitor-get-mounted-
volumes.

If the volume has an associated <gnome-vfs-drive>, it also appears in the drive’s list of mounted volumes, which can be queried using gnome-vfs-drive-get-mounted-

volumes.
[Signal on <gnome-vfs-volume-monitor>] volume-pre-unmount
   (arg0 <gnome-vfs-volume>)

This signal is emitted when the <gnome-vfs-volume>volume is about to be unmounted. When the volume is unmounted, it is removed from the volume-monitor’s list of mounted volumes, which can be queried using gnome-vfs-volume-monitor-get-

mounted-volumes.

If the volume has an associated <gnome-vfs-drive>, it is also removed from in the drive’s list of mounted volumes, which can be queried using gnome-vfs-drive-get-

mounted-volumes.

When a client application receives this signal, it must free all resources associated with the volume, for instance cancel all pending file operations on the volume, and cancel all pending file monitors using gnome-vfs-monitor-cancel.

[Signal on <gnome-vfs-volume-monitor>] volume-unmounted
   (arg0 <gnome-vfs-volume>)

This signal is emitted after the <gnome-vfs-volume>volume has been unmounted. When the volume is unmounted, it is removed from the volume-monitor’s list of mounted volumes, which can be queried using gnome-vfs-volume-monitor-get-

mounted-volumes.

If the volume has an associated <gnome-vfs-drive>, it is also removed from in the drive’s list of mounted volumes, which can be queried using gnome-vfs-drive-get-

mounted-volumes.

[Signal on <gnome-vfs-volume-monitor>] drive-connected
   (arg0 <gnome-vfs-drive>)

This signal is emitted when the <gnome-vfs-drive>drive has been connected. When the drive is connected, it is present in the volume-monitor’s list of connected

drives, which can be queried using gnome-vfs-volume-monitor-get-connected-
drives.
[Signal on <gnome-vfs-volume-monitor>] drive-disconnected
   (arg0 <gnome-vfs-drive>)

This signal is emitted after the <gnome-vfs-drive>drive has been disconnected. When the drive is disconnected, it is removed from the volume-monitor’s list of connected drives, which can be queried using gnome-vfs-volume-monitor-get-

connected-drives.
[Function] gnome-vfs-get-volume-monitor
   ⇒ (ret <gnome-vfs-volume-monitor>)
Returns a pointer to the <gnome-vfs-volume-monitor> singleton. <gnome-vfs-

volume-monitor> is a singleton, this means it is guaranteed to exist and be valid until gnome-vfs-shutdown is called. Consequently, it doesn’t need to be refcounted since gnome-vfs will hold a reference to it until it is shut down.

ret  a pointer to the <gnome-vfs-volume-monitor> singleton.

Since 2.6

13 MIME typing

functions to get a mime-type for a file using its name or its content

13.1 Overview

13.2 Usage

[Function] gnome-vfs-mime-type-from-name (filename mchars) ⇒ (ret mchars)

‘gnome_vfs_mime_type_from_name’ is deprecated and should not be used in newlywritten code. This function is deprecated, use gnome-vfs-get-mime-type-for-name instead.

Determine the mime type for filename.

Returns:

filename a filename (the file does not necessarily exist).

ret the mime-type for this filename. Will return ‘GNOME_VFS_MIME_TYPE_UNKNOWN’ if mime-type could not be found.

[Function] gnome-vfs-get-mime-type-common (uri <gnome-vfsuri>)
   ⇒ (ret mchars)

Tries to guess the mime type of the file represented by uri. Favors using the file data to the uri extension. Handles uri of a non-existent file by falling back on returning a type based on the extension. If cant find the mime-type based on the extension also then returns ’application/octet-stream’.

FIXME: This function will not necessarily return the same mime type as doing a get file info on the text uri.

uri a real file or a non-existent uri.

ret the mime-type for uri.

[Function] gnome-vfs-get-mime-type-from-uri (uri <gnome-vfsuri>)
   ⇒ (ret mchars)

Tries to guess the mime type of the file uri by checking the file name extension. Works on non-existent files.

uri a file uri.

ret the mime-type for file at uri.

[Function] gnome-vfs-mime-type-is-supertype (mime type mchars)

⇒ (ret bool) mime-type

a const char * identifying a mime type.

ret Whether mime-type is of the form "foo‘/’*".

[Function] gnome-vfs-mime-info-cache-reload (dir mchars)

Reload the mime information for the dir.

dir directory path which needs reloading.

[Function] gnome-vfs-mime-reload

Reload the MIME database.

[Function] gnome-vfs-mime-shutdown

‘gnome_vfs_mime_shutdown’ is deprecated and should not be used in newly-written code. This function doesn’t have to be called as the operating system automatically cleans up resources when exiting. Unload the MIME database from memory.

14 gnome-vfs-mime-utils

14.1 Overview

14.2 Usage

[Function] gnome-vfs-mime-type-is-equal (a mchars) (b mchars) ⇒ (ret bool)

Compares two mime types to determine if they are equivalent. They are equivalent if and only if they refer to the same mime type.

a a const char * containing a mime type, e.g. "image/png".

b a const char * containing a mime type, e.g. "image/png".

ret ‘#t’, if a and b are equivalent mime types.

[Function] gnome-vfs-get-mime-type-for-name (filename mchars)
   ⇒ (ret mchars)

Determine the mime type for filename. The file filename may not exist, this function does not access the actual file. If the mime-type cannot be determined,

‘GNOME_VFS_MIME_TYPE_UNKNOWN’ is returned.

filename a filename.

ret the mime-type for this filename or ‘GNOME_VFS_MIME_TYPE_UNKNOWN’ if mime-type could not be determined.

Since 2.14

[Function] gnome-vfs-get-mime-type (text uri mchars) ⇒ (ret mchars)

Determine the mime type of text-uri. The mime type is determined in the same way as by gnome-vfs-get-file-info. This is meant as a convenient function for times when you only want the mime type.

text-uri path of the file for which to get the mime type.

ret The mime type, or ‘#f’ if there is an error reading the file.

[Function] gnome-vfs-get-slow-mime-type (text uri mchars) ⇒ (ret mchars)

Determine the mime type of text-uri. The mime type is determined in the same way as by gnome-vfs-get-file-info. This is meant as a convenience function for times when you only want the mime type.

text-uri URI of the file for which to get the mime type

ret The mime type, or NULL if there is an error reading the file.

Since 2.14

15 MIME Database

functions for getting information about applications and components associated with MIME types.

15.1 Overview

This API can be used to query the applications and components associated with particular MIME types, and to collect extra information about them, and also provides a convenience function for launching them. Applications can register themselves for particular MIME types by adhering to the Freedesktop.Org Desktop Entry Specification

(http://standards.freedesktop.org/desktop-entry-spec/latest).

15.2 Usage

[Class] <gnome-vfs-mime-application>

Opaque pointer. This class defines no direct slots.

[Function] gnome-vfs-mime-get-all-applications (mime type mchars)
   ⇒ (ret glist-of)

Return an alphabetically sorted list of <gnome-vfs-mime-application> data structures representing all applications in the MIME database registered to handle files of MIME type mime-type (and supertypes).

mime-type a const char * containing a mime type, e.g. "image/png".

ret  a <g-list> * where the elements are <gnome-vfs-mime-application>

* representing applications that handle MIME type mime-type.

[Function] gnome-vfs-mime-get-all-components (mime type mchars)
   ⇒ (ret glist-of)

‘gnome_vfs_mime_get_all_components’ is deprecated and should not be used in newly-written code.

mime-type ret

[Function] gnome-vfs-mime-get-description (mime type mchars)
   ⇒ (ret mchars)

Query the MIME database for a description of the mime-type.

mime-type the mime type.

ret description of MIME type mime-type.

[Function] gnome-vfs-mime-can-be-executable (mime type mchars)

⇒ (ret bool) Check whether files of mime-type might conceivably be executable. Default for known types if ‘#f’. Default for unknown types is ‘#t’. mime-type

a const char * containing a mime type.

ret ‘#t’ if files of mime-type can be executable, ‘#f’ otherwise.

[Function] gnome-vfs-mime-application-launch
   (self <gnome-vfs-mime-application>) (uris glist-of)
   ⇒ (ret <gnome-vfs-result>)

Launches the given mime application with the given parameters. Command line parameters will be expanded as required by the application. The application will also be launched in a terminal if that is required. If the application only supports one argument per instance then multiple instances of the application will be launched.

app  the <gnome-vfs-mime-application> to launch.
uris  parameters for the <gnome-vfs-mime-application>.
ret  ‘GNOME_VFS_OK’ if the application was launched. ‘GNOME_VFS_ERROR_NOT_SUPPORTED’

if the uri protocol is not supported by the application. ‘GNOME_VFS_ERROR_PARSE’ if the application command can not be parsed. ‘GNOME_VFS_ERROR_LAUNCH’ if the application command can not be launched. ‘GNOME_VFS_ERROR_INTERNAL’ for other internal and GConf errors.

Since 2.4

[Function] gnome-vfs-mime-application-get-name
   (self <gnome-vfs-mime-application>) ⇒ (ret mchars)

Returns the name of the application app

app  a <gnome-vfs-mime-application>.

ret the name of the application.

Since 2.10

[Function] gnome-vfs-mime-application-get-icon
   (self <gnome-vfs-mime-application>) ⇒ (ret mchars)

Returns an icon representing the specified application.

app  a <gnome-vfs-mime-application>.

ret the filename of the icon usually without path information, e.g. "gediticon.png", and sometimes does not have an extension, e.g. "gnome-pdf" if the icon is supposed to be image type agnostic between icon themes. Icons are generic, and not theme specific.

Since 2.10

[Function] gnome-vfs-mime-application-get-exec
   (self <gnome-vfs-mime-application>) ⇒ (ret mchars)

Returns the program to execute, possibly with arguments and parameter variables, as specified by the Desktop Entry Specification

(http://standards.freedesktop.org/desktop-entry-spec/).

app  a <gnome-vfs-mime-application>.

ret the command line to execute. Since 2.10

[Function] gnome-vfs-mime-application-equal
   (self <gnome-vfs-mime-application>)
   (app b <gnome-vfs-mime-application>) ⇒ (ret bool)

Compare app-a and app-b.

app-a  a <gnome-vfs-mime-application>.
app-b  a <gnome-vfs-mime-application>.

ret ‘#t’ if app-a and app-b are equal, ‘#f’ otherwise. Since 2.10

16 Undocumented

The following symbols, if any, have not been properly documented.

16.1 (gnome gw gnome-vfs)

[Variable] gnome-vfs-ace-add-perm
[Variable] gnome-vfs-ace-check-perm
[Variable] gnome-vfs-ace-copy-perms
[Variable] gnome-vfs-ace-del-perm
[Variable] gnome-vfs-ace-equal
[Variable] gnome-vfs-ace-get-id
[Variable] gnome-vfs-ace-get-inherit
[Variable] gnome-vfs-ace-get-kind
[Variable] gnome-vfs-ace-get-negative
[Variable] gnome-vfs-ace-set-id
[Variable] gnome-vfs-ace-set-inherit
[Variable] gnome-vfs-ace-set-kind
[Variable] gnome-vfs-ace-set-negative
[Variable] gnome-vfs-acl-clear
[Function] gnome-vfs-acl-get-ace-list
[Variable] gnome-vfs-acl-kind-to-string
[Variable] gnome-vfs-acl-new
[Variable] gnome-vfs-acl-perm-to-string
[Variable] gnome-vfs-acl-set
[Variable] gnome-vfs-acl-unset
[Variable] gnome-vfs-context-check-cancellation-current
[Variable] gnome-vfs-drive-needs-eject
[Variable] gnome-vfs-escape-host-and-path-string
[Variable] gnome-vfs-format-file-size-for-display
[Function] gnome-vfs-get-default-browse-domains
[Variable] gnome-vfs-get-mime-type-from-file-data
[Variable] gnome-vfs-get-supertype-from-mime-type
[Variable] gnome-vfs-init
[Variable] gnome-vfs-initialized

[Variable] gnome-vfs-is-executable-command-string
[Variable] gnome-vfs-make-uri-canonical-strip-fragment
[Variable] gnome-vfs-make-uri-from-input-with-dirs
[Variable] gnome-vfs-make-uri-from-input-with-trailing-ws
[Variable] gnome-vfs-mime-application-get-binary-name
[Variable] gnome-vfs-mime-application-get-desktop-file-path
[Variable] gnome-vfs-mime-application-get-desktop-id
[Variable] gnome-vfs-mime-application-get-generic-name
[Variable] gnome-vfs-mime-application-get-startup-wm-class
[Variable] gnome-vfs-mime-application-new-from-desktop-id
[Variable] gnome-vfs-mime-application-requires-terminal
[Variable] gnome-vfs-mime-application-supports-startup-notification
[Variable] gnome-vfs-mime-application-supports-uris
[Function] gnome-vfs-mime-get-all-applications-for-uri
[Variable] gnome-vfs-mime-get-default-application
[Variable] gnome-vfs-mime-get-default-application-for-uri
[Variable] gnome-vfs-mime-type-from-name-or-default
[Variable] gnome-vfs-mime-type-get-equivalence
[Variable] gnome-vfs-module-callback-pop
[Variable] gnome-vfs-shutdown
[Variable] gnome-vfs-unescape-string-for-display
[Variable] gnome-vfs-uri-extract-short-path-name
[Variable] gnome-vfs-uri-get-fragment-identifier
[Variable] gnome-vfs-uri-make-full-from-relative
[Variable] gnome-vfs-volume-get-filesystem-type
[Function] gnome-vfs-volume-monitor-get-connected-drives
[Variable] gnome-vfs-volume-monitor-get-drive-by-id
[Function] gnome-vfs-volume-monitor-get-mounted-volumes
[Variable] gnome-vfs-volume-monitor-get-volume-by-id
[Variable] gnome-vfs-volume-monitor-get-volume-for-path

Type Index 40

Type Index

<gnome-vfs-drive>. . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 <gnome-vfs-file-info> . . . . . . . . . . . . . . . . . . . . . . . 13

<gnome-vfs-mime-application> . . . . . . . . . . . . . . . 35

                                                <gnome-vfs-volume-monitor> . . . . . . . . . . . . . . . . . 30

<gnome-vfs-volume> . . . . . . . . . . . . . . . . . . . . . . . . . . 21
<gnome-vfsuri> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Function Index 41

Function Index

C

compare . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21, 26

D

drive-connected on
<gnome-vfs-volume-monitor>. . . . . . . . . . . . . 31
drive-disconnected on
<gnome-vfs-volume-monitor>. . . . . . . . . . . . . 31

G

get-activation-uri . . . . . . . . . . . . . . . . . . . . . . . 22, 27 get-device-path . . . . . . . . . . . . . . . . . . . . . . . . . . 22, 27 get-device-type . . . . . . . . . . . . . . . . . . . . . . . . . . 22, 28 get-display-name . . . . . . . . . . . . . . . . . . . . . . . . . 22, 28 get-drive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 get-hal-udi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23, 28 get-icon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23, 28 get-id. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23, 28 get-mounted-volumes . . . . . . . . . . . . . . . . . . . . . . . . . 29 get-volume-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

gnome-vfs-acl-get-ace-list . . . . . . . . . . . . . . . . . 38
gnome-vfs-check-same-fs. . . . . . . . . . . . . . . . . . . . . 16
gnome-vfs-check-same-fs-uris . . . . . . . . . . . . . . . 16
gnome-vfs-connect-to-server . . . . . . . . . . . . . . . . 25

gnome-vfs-create. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

gnome-vfs-create-symbolic-link. . . . . . . . . . . . . 17

gnome-vfs-create-uri . . . . . . . . . . . . . . . . . . . . . . . . 14

gnome-vfs-drive-compare. . . . . . . . . . . . . . . . . . . . . 26
gnome-vfs-drive-get-activation-uri . . . . . . . . 27
gnome-vfs-drive-get-device-path . . . . . . . . . . . 27
gnome-vfs-drive-get-device-type . . . . . . . . . . . 28
gnome-vfs-drive-get-display-name . . . . . . . . . . 28
gnome-vfs-drive-get-hal-udi . . . . . . . . . . . . . . . . 28
gnome-vfs-drive-get-icon . . . . . . . . . . . . . . . . . . . 28

gnome-vfs-drive-get-id . . . . . . . . . . . . . . . . . . . . . . 28

gnome-vfs-drive-get-mounted-volumes . . . . . . . 29
gnome-vfs-drive-is-connected . . . . . . . . . . . . . . . 29
gnome-vfs-drive-is-mounted . . . . . . . . . . . . . . . . . 29
gnome-vfs-drive-is-user-visible . . . . . . . . . . . 29
gnome-vfs-escape-path-string . . . . . . . . . . . . . . . . 9

gnome-vfs-escape-set . . . . . . . . . . . . . . . . . . . . . . . . 10

gnome-vfs-escape-slashes . . . . . . . . . . . . . . . . . . . 10
gnome-vfs-escape-string . . . . . . . . . . . . . . . . . . . . . . 9
gnome-vfs-expand-initial-tilde. . . . . . . . . . . . . 11
gnome-vfs-format-uri-for-display . . . . . . . . . . . 9
gnome-vfs-get-default-browse-domains. . . . . . 38
gnome-vfs-get-file-info. . . . . . . . . . . . . . . . . . . . . 19
gnome-vfs-get-local-path-from-uri . . . . . . . . . 11
gnome-vfs-get-mime-type. . . . . . . . . . . . . . . . . . . . . 34
gnome-vfs-get-mime-type-common. . . . . . . . . . . . . 32
gnome-vfs-get-mime-type-for-name . . . . . . . . . . 34
gnome-vfs-get-mime-type-from-uri . . . . . . . . . . 32

gnome-vfs-get-slow-mime-type . . . . . . . . . . . . . . . 34
gnome-vfs-get-uri-from-local-path . . . . . . . . . 11
gnome-vfs-get-uri-scheme . . . . . . . . . . . . . . . . . . . 12
gnome-vfs-get-volume-monitor . . . . . . . . . . . . . . . 31
gnome-vfs-icon-path-from-filename . . . . . . . . . 12
gnome-vfs-is-primary-thread . . . . . . . . . . . . . . . . 12
gnome-vfs-make-directory . . . . . . . . . . . . . . . . . . . 20
gnome-vfs-make-directory-for-uri . . . . . . . . . . 20
gnome-vfs-make-path-name-canonical . . . . . . . . 10
gnome-vfs-make-uri-canonical . . . . . . . . . . . . . . . 10
gnome-vfs-make-uri-from-input . . . . . . . . . . . . . . 11
gnome-vfs-make-uri-from-shell-arg . . . . . . . . . 11
gnome-vfs-mime-application-equal . . . . . . . . . . 37
gnome-vfs-mime-application-get-exec . . . . . . . 36
gnome-vfs-mime-application-get-icon . . . . . . . 36
gnome-vfs-mime-application-get-name . . . . . . . 36
gnome-vfs-mime-application-launch . . . . . . . . . 36
gnome-vfs-mime-can-be-executable . . . . . . . . . . 35
gnome-vfs-mime-get-all-applications . . . . . . . 35
gnome-vfs-mime-get-all-applications-for-uri

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

gnome-vfs-mime-get-all-components . . . . . . . . . 35
gnome-vfs-mime-get-description. . . . . . . . . . . . . 35
gnome-vfs-mime-info-cache-reload . . . . . . . . . . 33

gnome-vfs-mime-reload . . . . . . . . . . . . . . . . . . . . . . . 33

gnome-vfs-mime-shutdown. . . . . . . . . . . . . . . . . . . . . 33
gnome-vfs-mime-type-from-name . . . . . . . . . . . . . . 32
gnome-vfs-mime-type-is-equal . . . . . . . . . . . . . . . 34
gnome-vfs-mime-type-is-supertype . . . . . . . . . . 32

gnome-vfs-move . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gnome-vfs-move-uri . . . . . . . . . . . . . . . . . . . . . . . . . . 15 gnome-vfs-open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 gnome-vfs-open-uri . . . . . . . . . . . . . . . . . . . . . . . . . . 15

gnome-vfs-remove-directory . . . . . . . . . . . . . . . . . 20
gnome-vfs-remove-directory-from-uri . . . . . . . 20
gnome-vfs-result-from-errno . . . . . . . . . . . . . . . . . 2
gnome-vfs-result-from-errno-code . . . . . . . . . . . 2
gnome-vfs-result-from-h-errno . . . . . . . . . . . . . . . 2
gnome-vfs-result-from-h-errno-val . . . . . . . . . . 2
gnome-vfs-result-to-string . . . . . . . . . . . . . . . . . . 2
gnome-vfs-set-file-info. . . . . . . . . . . . . . . . . . . . . 19

gnome-vfs-truncate . . . . . . . . . . . . . . . . . . . . . . . . . . 18 gnome-vfs-truncate-uri . . . . . . . . . . . . . . . . . . . . . . 18

gnome-vfs-unescape-string . . . . . . . . . . . . . . . . . . 10

gnome-vfs-unlink. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

gnome-vfs-unlink-from-uri . . . . . . . . . . . . . . . . . . 15
gnome-vfs-uri-append-file-name. . . . . . . . . . . . . . 4
gnome-vfs-uri-append-path . . . . . . . . . . . . . . . . . . . 4
gnome-vfs-uri-append-string . . . . . . . . . . . . . . . . . 4

gnome-vfs-uri-dup. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 gnome-vfs-uri-equal . . . . . . . . . . . . . . . . . . . . . . . . . . 7 gnome-vfs-uri-exists . . . . . . . . . . . . . . . . . . . . . . . . 17

gnome-vfs-uri-extract-dirname . . . . . . . . . . . . . . . 7
gnome-vfs-uri-extract-short-name . . . . . . . . . . . 8
gnome-vfs-uri-get-host-name . . . . . . . . . . . . . . . . . 5

Function Index 42

gnome-vfs-uri-get-host-port . . . . . . . . . . . . . . . . . 6
gnome-vfs-uri-get-parent. . . . . . . . . . . . . . . . . . . . . 5
gnome-vfs-uri-get-password . . . . . . . . . . . . . . . . . . 6

gnome-vfs-uri-get-path . . . . . . . . . . . . . . . . . . . . . . . 7

gnome-vfs-uri-get-scheme. . . . . . . . . . . . . . . . . . . . . 6
gnome-vfs-uri-get-user-name . . . . . . . . . . . . . . . . . 6
gnome-vfs-uri-has-parent. . . . . . . . . . . . . . . . . . . . . 5

gnome-vfs-uri-is-local . . . . . . . . . . . . . . . . . . . . . . . 5

gnome-vfs-uri-is-parent . . . . . . . . . . . . . . . . . . . . . . 7
gnome-vfs-uri-list-parse. . . . . . . . . . . . . . . . . . . . . 8

gnome-vfs-uri-new. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

gnome-vfs-uri-resolve-relative. . . . . . . . . . . . . . 3
gnome-vfs-uri-resolve-symbolic-link . . . . . . . . 4
gnome-vfs-uri-set-host-name . . . . . . . . . . . . . . . . . 6
gnome-vfs-uri-set-host-port . . . . . . . . . . . . . . . . . 6
gnome-vfs-uri-set-password . . . . . . . . . . . . . . . . . . 7
gnome-vfs-uri-set-user-name . . . . . . . . . . . . . . . . . 6
gnome-vfs-uri-to-string . . . . . . . . . . . . . . . . . . . . . . 5

gnome-vfs-uris-match . . . . . . . . . . . . . . . . . . . . . . . . 12 gnome-vfs-url-show . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

gnome-vfs-volume-compare . . . . . . . . . . . . . . . . . . . 21
gnome-vfs-volume-get-activation-uri . . . . . . . 22
gnome-vfs-volume-get-device-path . . . . . . . . . . 22
gnome-vfs-volume-get-device-type . . . . . . . . . . 22
gnome-vfs-volume-get-display-name . . . . . . . . . 22
gnome-vfs-volume-get-drive . . . . . . . . . . . . . . . . . 23
gnome-vfs-volume-get-hal-udi . . . . . . . . . . . . . . . 23
gnome-vfs-volume-get-icon . . . . . . . . . . . . . . . . . . 23
gnome-vfs-volume-get-id. . . . . . . . . . . . . . . . . . . . . 23
gnome-vfs-volume-get-volume-type . . . . . . . . . . 23
gnome-vfs-volume-handles-trash. . . . . . . . . . . . . 24

                                                gnome-vfs-volume-is-mounted . . . . . . . . . . . . . . . . 24
                                                gnome-vfs-volume-is-read-only . . . . . . . . . . . . . . 24
                                                gnome-vfs-volume-is-user-visible . . . . . . . . . . 24
                                                gnome-vfs-volume-monitor-get-connected-

drives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

                                                gnome-vfs-volume-monitor-get-mounted-

volumes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

H

handles-trash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

I

is-connected . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 is-mounted . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24, 29 is-read-only . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 is-user-visible . . . . . . . . . . . . . . . . . . . . . . . . . . 24, 29

V

                                                volume-mounted on <gnome-vfs-drive>. . . . . . . . 26
                                                volume-mounted on <gnome-vfs-volume-monitor>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

                                                volume-pre-unmount on <gnome-vfs-drive> . . . 26
                                                volume-pre-unmount on
                                                    <gnome-vfs-volume-monitor>. . . . . . . . . . . . . 30
                                                volume-unmounted on <gnome-vfs-drive> . . . . . 26
                                                volume-unmounted on
                                                    <gnome-vfs-volume-monitor>. . . . . . . . . . . . . 30